You should be able to solve this
>>689912733
50%
>>689912733
1 crit / 2 crit / no crit
33.3 %
>>689912733
75%
>>689913269
but 1 is 100% crit
>>689913269
>no crit
>at least one of them is a crit
Wew lad
its 50%
because you hit him twice.One of the hits are already known to be a crit for both of them to crit you just need to crit once meaning its 50%
>>689912733
question is made by someone who thinks he is a mathematical genius or something but in reality the question is fucking retarded
>>689912733
1/2 * 1/2 = 1/4 or 25% chance
>>689913404
>>689913426
oh then what >>689913103 said /thread
pack it up boys shows over
ATLEAST ONE HIT IS A CRIT, NO MATTER FUCKING WHAT. NOW WE ONLY HAVE THE OTHER HIT TO LOOK AT. SO 50%
1: 1st crit, 2nd crit
2: 1st crit, 2nd don't crit
3: 1st don't crit, 2nd crit
1/3 chance.
>>689913269
>At least one of the hits is a crit
First hit= crit 100%
>Assuming 50% crit chance
Second hit= crit 50%/no crit 50%
= 50%
>>689912733
50%, suck my dick, nigga
>>689912733
It's literally crit or not crit one of the times, so you can eliminate that time. After that, it's just 50/50
So much ameritard here.... 0.25 is the answer
>>689913746
this anon gets it
>>689912733
Mathematican here,
Let A be the first hit and B the second. There are 4 possibilities:
A=crit B=crit
A=crit B=x
A=x B=crit
A=x B=x
We have the condition that at least one hit is crit so we only consider the first three cases. That means the possibility is
1/3= 33.333%
>>689912733
75%
100%(0.5)+50%(0.5)=75%
>>689912733
1 in 2
1 crit + depend on how many probabilities?
0.25
>>689913269
Right answer, wrong method. Draw a punnett square: there are 4 possible outcomes (ab, Ab, aB, AB). We know from the question that one of them (ab) is off the table and our solution is one of the other three. Of the other three, what percentage are double-crit? 33%.
4 different possible outcomes. 2 of them have 2 criticals. 50% chance.
This is the same as the coin problem. It's 1/3.
>>689912733
66%
can someone please legit answer this
â…“ you idiots
>>689913883
>at least one of them is a crit
>A= x B= x
pure autism
>>689912733
82%
>>689913883 A=x and B=x can't happen. And if you are counting all possible outcomes where we know one is critical there are 4 possible outcomes with 2 of them having 2 criticals.
>>689912733
It's obviously 50%. Next question.
>>689914258
its 50% mate
>>689914120
Yep easy. Make the problem harder. You hit the enemy 4 times the 1st and 3rd hits are crits. You got at least 3 crits. What is the probability all hits were crits?
>>689912733
73%
>>689912733
37.5% chance
>>689912733
1/2 * 1/2 = 1/4.
C = Crit, N = No crit
NN, CN, NC, CC
Only one instance of four possible outcomes is both crits.
25% chance.
67%
.3 repeating.
>35 replied
>25 posters
fuck off
>>689914334
>>689914394
>We have the condition that at least one hit is crit so we only consider the first three cases.
Do you autists even read?
>>689913833
the question is "what's the probality BOTH hits are crits?"
C:0.5 then C again=0.5 (or /C=0.5) so 0.5*0.5= 25%....
>>at least one of the hits is a crit
So fcking useless cause the question is BOTH are crits
fags
Still 50% assuming the hits are independent of each other. With one of the two trials having already been completed, we can look at the other hit isolated.
>>689914394
>A=x and B=x can't happen
read my post again
>here are 4 possible outcomes with 2 of them having 2 criticals.
No. If you don't understand probability theory than don't post sit.
If someone's interested: This problem is called the "Boy or Girl paradox", see
https://en.wikipedia.org/wiki/Boy_or_Girl_paradox
>>689914642
Except there aren't 4 instances because one is a guaranteed crit dumbass.
>>689914871 did you read the rest of my post. Please explain.
Mathematician here: 50%
Every scientific percentile is a 50/50 chance when devised by retards like you. Either the desired effect occurs or it does not.
>>689915133
No, see >>689913883 or the link in >>689915002
Your fault is that you assumed that A and B are independent. That's not the case.
P(crit n crit | crit) = (crit n crit n crit) / crit = (0.5 × 0.5 × 0.5) / 0.5 = 0.25
>>689912733
1/3 or 33% no arguing.
>>689915133
Wrong. You are not a mathematician.
>>689912733
50....
>>689912733
HC
CH
CC
All these are equally likely to happen since the chance to crit each time is 50%.
The probability of hitting both crits=1/3
Every other answer is wrong.
>>689912733
50 percent u retard
>>689913670
That's not how it works.
>>689913810
1/3 nigga
>>689915318
forget to write text...
the chance that both are crits is 0,25
we can't just ignore the possibility that both won't crit
>>689913833
That's very far from the answer. It's 1/3
the probability of the pool is not affected by the base chance. so the probability that the other hit is crit is just 50%. it isnt affected by the other one.
>>689913883
Yes but you sound like an asshole when you say mathematician here.
>>689912733
This >>689913103 all things in life boil down to it will or it won't.
>>689914258
1/3
>>689912733
there are only 3 cases possible:
First hit crit, the second don't crit.
First hit crit, second crit
First hit don't crit, the second crit.
so there are 33% chance of a double crit.
>>689913567
the correct answer
>>689915807
We can, it's literally stated in the problem that at least one of the attacks will crit.
>>689915680
a voice of reason in a sea of filth
>>689915807
Yes we can. That's what called "conditional probability".
>>689912733
well you calculate each one indenepndently, its not like well the first one was a crit so 1 is 50% of 2 and it was a crit so 2 wont be a crit..
no its each on its own, at each hit its 50-50 crit or not. In theory you could hit 1000 times and non of them could be a crit,, or all of them could be, if you were lucky. Altho 50% is pretty large so thats unlikely but,, not impossible.
so one is a crit no matter what, then its 50%, since its either the other is a crit or not, no other option.
>>689915807
Yes we can. Yes we do.
>>689916115
Dear God...
The world is heading into dark places.
>>689915425
I object.
The proof is kinda complicated, but regarding the theory of conditional theoretical chances, the answer is 0.5
>>689912733
50%
since the crit chance doesnt drop after a crit is made each attack has an independent 50% to crit
>>689912733
25% chance. ezpz.
>>689915995
Well, at the moment I'm working on my master thesis in math so I would say I'm a mathematician.
>>689916251
thats my thought as well. the second hit isnt dependent on the first and the result of the first hit has already been established.
>mfw niggers really arguing the order in which the crit is going to occur and making it a probability in outcome percentage
its 50%
>>689916108
Thank this guy
>https://www.youtube.com/user/MindYourDecisions
His videos taught me a lot about probability and logic
>math is hard
Jesus Christ, it's fucking 33%, if you're really that stupid just use Google.
>>689916188
No you can't treat them as separate events one the question is about multiple events happening.
>>689916209
explain
The people who don't say 1/3 are the same people who don't know the difference between your and you're
itx 33%, any other answer is mentally fucking retarded and cant do simple math
>>689914258
25% obviously what the fuck is wrong with you dumb mongs
>>689916393
Fine you can look at it as 1 crit or two cits, just 1 crit is twice as likely so you know, still 33%
Anyone who says something other than 1/3 is clearly underage, as anyone who's completed high school should be able to calculate such basic conditional probabilities.
>>689916232
Wrong, it's 1/3. Rather famous question actually.
Three possible outcomes, one of them has two critical hits. Answer is therefore 1/3.
>>689914258
1/3ed
normally there are 4 outcomes
nc/nc
c/nc
nc/c
c/c
The question removes nc/nc as an answer
This leaves us with
c/nc
nc/c
c/c
so it's 1/3
A lot of peopel say 1/2 because they make the incorrect assumption that nc/c and c/nc are the same.
They are retarded.
>>689916393
Wrong, it's 1/3. Rather famous question actually.
Three possible outcomes, one of them has two critical hits. Answer is therefore 1/3.
On a table I have an apple and a coin. If the apple is an apple and the coin is a fair coin, what's the chance that a coin flip will result in heads and me also having an apple on the table?
50%
>>689916645
Wrong, it's 1/3. Rather famous question actually.
Three possible outcomes, one of them has two critical hits. Answer is therefore 1/3.
>>689916765
one result has already been abrogated by assuming that we already have a crit. so the answer is .5
>>689916749
ding ding
>>689916506
No, we do not fucking ignore the possibility of neither hit being a critical. If that was the case, then would we not also ignore the possibility of both hits being crits?
Say the question was 'what is the possibility both hits aren't crits?' Would that be impossible to work out?
>>689913567
This
>>689916796
Wrong, it's 1/3. Rather famous question actually.
Three possible outcomes, one of them has two critical hits. Answer is therefore 1/3.
>>689916907
you are wrong, but its ok to be retarded
>>689916341
everyone is just over complicating it when the question is literally whats the chance a hit will crit when its a 50-50 chance to crit.
theres no chance of missing since question never gives us a hit chance thus its guaranteed to hit
>>689913269
That's not how probabilities work :'(
>>689913883
So, if i wanna write algorithm for this shit, i must include your explanation?
This returns 25%:
import random
def rand():
return random.randrange(0,2) == 1
def test(count):
success = 0
for i in range(0, count):
if(rand()):
if(rand()):
success = success + 1
return (float(success) / count) * 100
print(test(10000), "%")
What wrong here?
>>689913746
yay 50%
>>689916907
Yes that's was I removed one which made it 3 instead of 4.
>>689916951
How can you get an outcome with no crits if one of the hits is critical?
>>689912733
If it is my teammate shooting, then 0% to hit anything.
>>689917035
Odly, it's the right answer.
P(crit + crit) = 1/2 * 1/2 = 1/4 = 25%
>>689916796
You have 2 coins on the table. Flip both coins until 2x heads, disregard any 2x tails. That leaves HT HH TH 1/3
33%
because "At least one of the hits is a crit."
>>689916933
This is this worst post I've seen all day.
50% fags are bad but 25%? Holy fuck.
>>689917065
1 critical has been established
then:
50% chance to crit
50% chance to not crit
so the chance of 1/1 is 50%
and the chance of 2/0 is 50%
>>689917081
I didn't include a scenario of both hits not being critical so the calculation is still accurate...
>>689916939
this allows for the event of 0 crits
>>689917285
>>689917279
Retarded
1/3
>>689912733
25%
>>689917279
No critical has been established, merely the case in which no critical has been ruled out.
no crit / crit
and
crit/ no crit are both still viable and still equally likely to crit/crit
>>689917279
You don't understand do you
It's 1/3 shut the fuck up you uneducated niggers
>>689917541
Wrong but its ok to be dumb :)
>>689917279
the crit hasn't been established, it could be the first or the second hit, making 1 crit twice as likely as 2
>>689917580
See it's not wrong at all.
>>689917037
You're a shit programmer you forgot to include the information 'at least one of has to be a crit' so it has to be
x = rand()
y = rand()
if(x == y and not (x==0 and y==0))
>>689917037
You're checking if both are cross, which is 25%
If you want to test the problem here then you only count results where at least 1 is a critical, then count how many of those qualifying results are both hit. Your brute force will return around 33%
>>689912733
Each hit has a 50% chance regardless of the number of hits you take, or the number of previous successes you have. 1 hit or 5 million, 50% is 50%.
>>689913883
>Mathematician here
>Autist here
>>689917632
whoops @@;
my bad yall, haha
>>689917644
oh it is, unintelligent human :)
>>689913567
Not at all.
One of the events is already a crit, so even by your flawed scheme it would be 1 * 1/2 = 1/2 or 50% chance.
Crit & Not Crit are mutually exclusive, so you ignore the probability of it not being a crit and instead work on the same basis as the whole 'bag full of coloured beads, take one out, don't put it back' scenario.
>>689914642
this
>>689915053
he's right tho... are you the one saying 75% cuz thats legit retarded. youre odds dont increase, they decrease. logic your way out of it dude.
>>689912733
1/3rd
>>689917279
youre assuming that one hit has been made and youre calculating the chance the next hit will be a crit. thats not the scenario in question.
>>689917721
No that is not how probability works.
>>689917788
I get it now, you're right. That's what I was doing
>>689917783
wrong
>>689917738
Please explain. I know you are wrong it's just fun to kick at people once they're down.
>>689917930
It's right.
>>689916933
Let me give you an example.
Someone tells you he's thinking about a number between 1 and 100.
What's the probability it's 1? Well it's 1% of course.
Now he also tells you his number is smaller than 3. What's now the probability his number is 1?
According to you it would still only be 1% but that's wrong. There are only two possibilitys: 1 or 2 so it's 50%. We can ignore all other numbers which not match the given condition.
Really, just look up condition probability.
>>689917764
You didn't expand the not critical tree. The second hit could be a critical, which would still fit into the statement "one of them is a critical"
>>689917935
dont need to, you wont get it anyways with your intelligence :-)
It's 0.25. What's the probability of you hitting ten (50%) crits in a row? 0.09765625%. Because you have to take into consideration the probability of every single crit that occurs.
>>689916657
how is it twice as likely when its either going to be
crit+crit or crit+non crit
>>689917788
1 is always a crit you mongoloid
>>689918084
This
>>689916933
>>689917037
It's given that there is at least one critical.
We all agree that there are four equally possible possibilities in two hits, unless you're retarded: two regular hits, first critical second regular, first regular second critical, both critical (RR, CR, RC, CC)
Now, it's given that there is AT LEAST one critical, so we're calculating P(X=2 | X>=1) or the probability that there are two criticals, given that there is at least one. Two critical hits fall within the given statement, therefore will be counted. The two chances of one critical fall within the given statement, therefore will be counted. The chance of no criticals, however, does not fall within the given statement, and will be excluded.
Now, we have, through the given statement, the possibilities (CR, RC, CC). Now, we're looking for the probability of (CC), which is one out of the three given possibilities. Therefore, the answer is 33%.
>>689918066
So I win if you can't disprove me?
>>689912733
>stop browsing /b/ for a few months
>come back and see this question with different wording for the 100th time
>people still think it's 50/50
never change /b/
>>689917935
You sound like the biggest edge lord believable
I was saying .5, it's clearly .33.
Just calm down and think about it.
>>689918205
no, i win because you actually think its 33%, but if you want to say you win for no reason, thats fine, its what children do :)
>>689916933
So you're saying what if the question was,
"what is the probability of getting no crits, but 1 hit is a crit?"
>>689918205
Don't take the bait fren
>>689918021
The question was 'what is the probability that BOTH hits are critical'. Your statement is irrelevant to the question at hand, we're not working out if 'one of them is critical'.
>>689918298
3 possible outcomes, one of them has two critical hits. 1/3.
1:4
>>689917714
>>689917693
Yeah, i forget about x == 0 and y ==0. Now it's getting near 33%. Now i understand.
>>689917731
My autist roommate failed his Probability Theory class specifically because of his inability to understand given statements.
>>689918122
flip 2 coins a few times, you'll quickly see that h/t happens about twice as often as h/h. Because in reality it's 2 different events one being h/t, the other being t/h. They just look the same so you can say it's one event that's just twice as likely
>>689913103
>>689913269
>>689913288
>>689913502
>>689913670
>>689913746
>>689913795
>>689913810
>>689913824
>>689913883
>>689913914
>>689914032
>>689914033
>>689914120
>>689914183
>>689914316
>>689914375
>>689914425
>>689914453
>>689914504
>>689914657
>>689915442
>>689915680
>>689915631
>>689915680
>>689915696
>>689915786
>>689915894
>>689916062
>>689916067
>>689916188
>>689916251
>>689916393
>>689916441
>>689916702
>>689916749
>>689916765
>>689916796
>>689916847
>>689916951
>>689917261
>>689917279
The "at least one of the hits is a crit" is superfluous information. If you had studied probability you'd know that the correct answer is 0.5*0.5 = 0.25 or 25%. All other answers are incorrect. Period.
>>689918416
you either hit or you dont, sorry friend, its 50/50 but good try, just work on the math a bit, ok? :)
>>689916095
This.
>>689918537
B8 right?
>>689918127
one of the two hits is a crit, not the first hit is a crit. not only do you fail at simple conditional probability, but your reading comprehension is quite poor too. dont worry champ, youll learn this by the time you graduate high school.
the acrimony in this thread is a good example of why people lose all their money gambling and making bad choices. probability isn't natural for most people. a lot, including myself, overestimated the chances by 17%
>>689918419
just FYI 1:4 means 20%:80% or 1/5:4/5 and not 1/4 or 25%
>>689918588
Just realized this is b8. :P
>>689918537
THe funny thing is the question is famous for having people argue between 1/3ed and 1/2, but you, you're a special little guy aren't you.
i just cant understand the level of retartidity in here
>>689918537
THIS. Please someone help these poor, imbecilic souls that don't understand this.
>>689917840
Yes it is, by giving the outcome of one of the hits it is no longer a question of probability. You're estimating the chance that the remaining hit is a crit, since the other has occurred. 1*.5 = .5 the probability is 25% before either hit is taken, its 50% after the first is confirmed.
>>689918537
Absolutely incorrect.
"At least one of them is crit" is essential information. It removes the possibility of them being both non-critical. So there's only 3 possibilities. CC, NC and CN.
Answer is 1/3 = 33%, not 25%.
>>689918537
>superfluous
Absolutely wrong. This information tells you to eliminate the fourth possibility and that's crucial to solving the problem.
>>689918537
>>689918610
if we take this bait will he start yelling about how he is a math genius and has a foot long dick that he fucks bitches with by impressing them with bad math?
>>689918740
samefag
>>689918610
No, moron. Look up how to solve probability problems if you wanna find out
>>689918769
Very famous question. You are arguing with Eeinsteins opinion just fyi.
>>689918619
I didn't say the first is always a crit. I said one of them is a crit.
>And your reading comprehension is quite poor too.
Nice irony
>>689918769
25% no crits
50% one crit
25% two crits
when no crits is removed, chance of two crits is 25/75 or 1/3
>>689918537
I checked it by script, statistics say, that you wrong.
>>689918740
You're both wrong.
I'll give you this information:
>At least one of my children is a boy.
Now answer this question:
>Are both my children girls?
>>689918878
Very famous question. You are arguing with Eeinsteins opinion just fyi.
>>689918675
I thought that's how you wrote probabilities too, no?
>>689918769
No hit already has an outcome, reread the question
"AT LEAST ONE OF X ARE Y" does not mean "THIS X IS Y" is just means in this group of Xes at least on is a Y.
>>689918694
Because it's 25%. It can't be a third, because there isn't three possibilities anyway. There's Crit-Not crit, Not crit- crit, crit- crit and not crit- not crit. Using you people's method (which is the way that 11 year olds do probability) shows that its 0.25. 1/2 x 1/2 = 1/4.
>>689918769
you're not given the outcome of a specific hit. you're simply told that the no crit scenario did not occur as at least one hit critted.
>>689918662
It's not so much why people lose gambling, but why people gamble at all. If you learn a decent amount of probability theory, you know that every casino and other gambling area is made so that the house should always have a significant profit, as long as the players aren't cheating or doing heavy cooperation.
>crit
>crit
>crit
>not crit
>not crit
>crit
>crit
>crit
possiblility of double crit is 2 out of 4 times
therefore 50% chance of having double crit
>>689918537
>The "at least one of the hits is a crit" is superfluous information
Are you actually brain dead? Having at least one as a crit eliminates the possibility of of a no crit outcome. The situation isn't "2 strikes with a 50% crit chance" it's "1 strike with 50% and the other is 100%"
Which leaves you with 3 possible outcomes not 4.
>>689919067
Very famous question. You are arguing with Eeinsteins opinion just fyi.
I have the ability to look at and immediately solve problems. The answer is 25%
>>689918694
>>689918740
>>689918828
>>689918847
>>689918859
Lol. I think the real problem here is that they didn't teach any of you enough math at school. All you gotta do is google real quick if you're really interested
>>689919116
I was assuming people gamble a little for fun. But some people think you can win at it, so they keep overestimating their chances and lose their shirt.
>>689919124
Literally retarded.
>>689919184
Wrong.
Very famous question. You are arguing against Eeinsteins opinion just fyi.
>>689919076
The question rules out not crit- not crit
gey your head checked familam
>>689919124
how do you even manage to breathe
>>689919017
it's how you write a ratio, if you wanted to write a probability like that you certainly could.
if you meant to say it had a 1/5 probability then it was the right way to put it, I just don't see how you could get that
>>689913914
I came up with the same thing.
Fucking retards. 0.5 × 0.5= 0.25 or 1/4 or 25%
>>689919222
Very famous question. You are arguing against Eeinsteins opinion just fyi. But I suppose he was wrong.
>>689919116
Actually if you know how to play you can actually have a higher probability to win than the house in Craps.
>>689919184
While it strongly depends on how the question is asked (and the OP question doesn't contain the necessary information to give an absolute solution) if the answer is 1/3 or 1/2, 25% is NEVER the correct solution.
>>689912733
Over all there's a 25% chance to crit two times
There's a 25 you don't crit at all
And 50 that you only crit once.
Since we know you hit once for x2dmg then we can go ahead and mark of the possibility of not landing any at all.
I'm not good with percentage but I know the fraction chance is
1/3
1 possibly for the first hit to be a crit
1 possibly for the second hit to crit
And 1 possibly for both hits to crit
It's a 33.3%.
Unless you're a stormtrooper.
Then it's 0%.
>>689919124
my sides
>>689919080
which means only one of the hits matters in the calculation, since the other is stated. it doesnt matter what order they're in, if you know what one of them is its not a question of probability anymore. you should only try to calculate the unknown, which has a 50%.
>>689912733
youve completely fucked up the question and those fucking retards are still fucking retards
>>689919124
66,6%
1. crit hit
2. crit crit
making the third 3. a possibility outcome that has nothing to do with the first 2 outcomes
>>689919222
Answer this post >>689918985 , then.
>>689918950
and yet you still cant get it right. never mind champ youll get there one day.
>>689919462
Fuck off traitor
>>689919421
i dont think thats true. and ive played craps quite a bit.
>>689919549
But how do you account for the difference in probability when the critical hit happens first and when the critical hit happens second?
>>689919400
Well if Einstein would have answered anything other than 25% to this question than yes he was wrong
>>689914928
This guy is the winner
>>689912733
4 equally likely outcomes:
CC
CN
NC
NN
Excluding NN from the outcome set as per the question, one out of three equally likely outcomes results in two crits, ergo:
1/3
>>689919678
yes, that is a good way of saying it
>>689919021
H = hit
N = no hit
Prob(NN) = 0%
If prob(HH) > 0% then prob (H) > 50%
Therefore the answer is 0%
>>689919258
I define "gambling for fun" as just playing, not gambling. When you do it for fun, you're not doing it for money, really.
Probably should've specified, though. Oh well.
gam·ble
ˈɡambəl/
verb
play games of chance for money; bet.
>>689912733
Easy, 75%
One is 100%, the other is a 50/50
>>689919716
I see. Do you have anyone backing you except for 4chan?
>>689912733
25%
>At least one is a crit
>1:1st hit crit, 2nd hit no crit
>2:1st hit no crit, 2nd hit crit
>3:1st hit crit, 2nd hit crit
1st hit no crit, 2nd hit no crit = DNE because "at least one is a crit"
>>689919754
But then the prob of a hit is 1/3 not 50%.
>>689919606
Get what right? Stop spending your time calling people 2 year olds and actually say something noteworthy. Unless the problem is that whenever you're wrong you just resort to insulting others.
>>689919614
Three things cannot be long hidden: the sun, the moon, and the truth - Buddha
Stormtroopers suck.
>>689914928
But that statement eliminates one outcome and is therefore essential to solving the problem. How do you not see that?
If that statement weren't in there your group would be
>C, noC
>noC, noC
>noC, C
>C, C
While WITH the statement it's
>noC, C
>C, noC
>C, C
That's a completely different issue.
>>689919803
Well, consider how risk itself is attractive for some people. They want the thrill of maybe losing their money more than they want the expected value of a hit. It's kind of semantics, but stuff to think about.
>>689919716
Except you're retarded
>>689912733
42%
>>689919716
25% is literally retarded.
I can see how intelligent people would get 50% and just be careless, but 25% means you are actually a stupid person.
>>689919716
can we all just take a moment to look at this.
>>689919888
I don't even know what you are trying to say but it's probably dumb.
>>689919549
incorrect. you have two hits, but you dont know which of them crit. could be just the first hit, could be just the second hit, or could be both. you cannot say that a specific hit critted, you do not have the information to confirm which hi crit, just that one or more did.
>>689919808
You're wrong, even by your own logic.
>>689919920
hes arguing against my point>>689916393
your answer is the correct one imo
>>689914928
>probability doesn't depend on the alternatives
where you even taught stats in school?
>crit - no crit
>no crit - crit
>crit - crit
>no crit - no crit
1 in 4
If you actually wen to school you would get 75%
>>689914334
You are the dumbest person I've met today, possibly this week so far.
>>689920224
he's thinking about the independence of each coin flip, he just doesn't understand how it applies in this case
25% no crit
50% 1 crit
25% both crit
>>689919421
I highly doubt that, but even if it's true, those who actually can control the game and themselves is so insignificant it doesn't bother the house. Proof? Look at this thread, maybe 0.01% of those who have the right answer can actually do that stuff, and probably 0.01% of those who can do it will actually put in the time and effort to try it. That's a really damn small number if you expand it to the entire population.
>>689912733
420%
>>689919678
Scenario 1
Hit 1 - CRIT
/ Not Crit
Hit 2 50%
\ Crit
Scenario 2
/ Not Crit
Hit 1 50%
\ Crit
Hit 2 CRIT
there are 2 possible scenarios and they both have a 50% chance.
>>689920166
Care to explain how?
One hit WILL be a crit. That is a 100% chance, or 1.
The other hit has a 50% chance, or .5, to crit.
1.5/2 = .75
or 75% chance to crit both times
>>689920332
>>689920262
Consider suicide for the sake of our gene pool.
33% both crit
66% 1 crit
>>689912733
Serious question, do you get taught probability and stats in US highschool?
>>689919888
What are you arguing here?
>>689916393
The order is important.
Suppose you throw a crit at your first hit, for the second crit you've now got 2 possibilities:
>no crit
>crit
Right? Good.
Now suppose your first hit is NOT a crit. In order for the initial condition to be true there is only one outcome left:
>crit
The order IS something to be factored in.
Generalized it looks like this:
>no crit -> crit
>crit -> no crit
>crit -> crit
That's the only possible scenarios and that's because the order is essential.
>>689920463
Even by your logic the answer would be 25% which still is very wrong
50% * 50%
it's 50% why do we even arguing
>>689915977
The question wasn't "your hit was a crit, what's the probability of your next one being a crit". The question was what's the chance of having 2 crits if at least one of them is a crit. Seems subtle but very different questions.
>>689912733
I've changed my mind. let's assume hit 1 doesn't crit. That would mean hit number 2 would have to crit. hit number 1 won't crit 50% of the time. hit number 1 will crit the other 50% of the time, which means hit #2 can either not crit, or crit. Since not critting both times is not a possible outcome it goes as follows.
H,C will happen 50% of the time, because the second hit will always crit in the event hit 1 does not based on the question.
CH will happen 25% of the time
CC will happen 25% of the time
You can flip coins and do it. every time it flips tails count the second flip as heads. every time it flips heads flip the second coin for heads or tails. It's 25%
>>689920463
1
first hit: crit
second hit: no crit
2
first hit: no crit
second hit: crit
3
first hit: crit
second hit: crit
4 (cant happen)
first hit: no crit
second hit: no crit
see this? i was missing it myself
>>689920478
SO close
>>689918537
One is guaranteed 100% to be critical. So. 1/1*1/2= 1/2 = 50%
>>689920472
>1.5/2
Spectacular retardation.
>>689920628
Because its 33%
>>689916188
You can't calculate each one independently, because the outcome of the first influences the possible outcomes of the second. If your first is a crit, your second can be either, if your first is not a crit, your second CAN ONLY be a crit. See >>689920569
[spoiler]>>689912733[/spoiler]
Wow OP, you're fucking stupid. obviously the chance is 500% + an active genital yeast infection divided by mitochondrial bacteria in its gestation phase.
>>689916749
My philosophy professor said the same thing.
>>689920683
you're the first one to explain it properly.
>>689918828
The option of CN is the same as NC. So there fore the only choices you really have are CN And CC. Thus 50%
>>689920463
the problem is, you don't know which bullet has the absolute crit so when the first bullet has crit you can wonder if that's pure chance or absolute.
>>689916749
sorry but most of these people haven't taken statistics. I'm assuming that your assumption about their assumption is incorrect. I assume that the place people struggle is with the sentence order. People assume that the first hit must be a crit because your given an event before the odds. In this case they believe the outcome will only be based on the odds of that last move
>>689920472
lol
no it's 25%
0.5 * 0.5 = 0.25 = 25%
there are 4 combinations
hit hit
hit crit
crit hit
crit crit
always 50/50 so 50% for the first being crit + 50% on the so it's 0.5^hits
>>689919978
We certainly are a wonderful species. From brain-dead retarded to Einstein smart, there are endless wonders and surprises to be found.
...but I still can't believe it's not butter.
>>689921000
Bruh hit hit cant happen.
>>689921000
sorry didn't really read the topic I calculated 50% crit chance
but still trips and dubs
Ahem.
So there's a 1/2 chance for a crit, and you want there are two possible outcomes. This means that having a crit happen the first time is a 50/50, and then having it hit the next time is also a 50/50. Considering that it happens twice, and the chances are both 1/2, theoretical probability tells us that we multiply the odds together.
50%×50%
|
50/100×50/100
|
1/2×1/2
Probability of a crit= 1/4
Remember that this doesn't even matter because it is a probability and therefore van have multiple outcomes, but this just describes the probability of this occuring.
25% only if you don't know the first is a crit. so it's 100%*50%=50% (first is 100% because we know for sure it lands a crit)
>>689920960
They are not the same.
Why would you think they are the same.
>>689920999
I never took stat and I still got it right because I'm not a retard.
>>689920453
This
>>689920472
>Care to explain how
By your own logic:
You assume your first crit is a hit. That means you only have to look at the second one, as the previous outcome is already known and the probability doesn't ever change, making it wholely depend on the probability of the second throw, which is 50%. By your reasoning you should have arrived at the 50% chance, not the 75%.
Real logic:
The order is something to be factored in and the fact that you don't know what hit is a crit influences the outcome as outlined here >>689920569
>>689920927
except he's wrong. you can't just remove no crit / no crit from the equation because if you don't crit on the first shot you will always crit on the second based on the question. That gives no crit / crit a 50% chance of happening, crit/nocrit 25% and crit/crit 25%
>>689915753
that is exactly how it works. given that one is already a crit we can ignore it and focus on the other one,which we know has a 50% chance of being a crit. therefore the chance that both are crits, given that one is going to be a crit, is 50%.
>>689921261
Why can't you remove it? That was Einsteins calculation so I don't see you winning this one.
>>689917764
Nothing has been taken out yet. The question wasn't "crit chance is 50%, and you already have a crit, what's the likelihood of your next being a crit" . If it was you'd be right, but that wasn't the question.
>>689920706
"one of them", yes. So it could be that the first one is critical, that the second one is critical, or that they're both critical. So there's 3 possibilities, not 2.
Answer is 1/3 = 33.3..%
>>689920292
What kinda fuckin school did you go to
>>689916095
it has been stipulated that the event of no crits is not part of the data we are collecting
>>689921261
if i crit on the first, i can normal or crit on the second. thats two outcomes.
if i normal on the first, i crit on the second. thats one outcomes
thats still one outcome with two crits out of three possible outcomes. yes?
>33%
IQ > 100
>50%
IQ 85 - 100
>25%
IQ < 85
>Anything Else
I'm sorry for your family
>>689921178
Because the order of the hits do not matter. One hit is 100% going to be critical. The other hit you take boils down to a 50/50 option. 1*1/2=1/2
For all the faggots who say it's 50% because they don't understand the concept of 'at least one of the hits is a crit'.
Imagine the following scenario:
First hit ends up not being a crit. You have already lost, BUT you can guarantee the next one will be a crit.
Only when the first one does and up hitting a crit, you have a 50% of getting the second crit. Out of these 3 just as likely scenarios, only one is the winning scenario (2 crits), which means you have in total a 1/3 chance of getting that scenario.
>>689920472
Figuratively retarded.
>take 1st grade math
>learn that .5 * .5 is 0.25
>take stats
>learn that .5 * .5 is 0.25
>american education system
>>689921465
Junior Negro College for cotton picking.
>>689918127
It always makes me laugh when dumb people think they're right and call others dumb. The irony is too much.
threads about to slide off
thanks for the conversation fellows
*tips fedora*
seriously though, it was fun to see all the explanations and answers
>>689919353
Oh right, ratios. No I meant 1 out of 4
>>689921535
The order does matter. At least that's what Einstein said.
>>689921402
Because the problem says YOU WILL ALWAYS CRIT ONE HIT. which means NOCRIT/NOCRIT is not possible. that means with 50 percent crit chance what are the odds your first hit won't crit? 50% which means your second shot will crit. That means 50% of the time its NOCRIT/CRIT. 25% of the time its CRIT/NOCRIT. 25% of the time its CRIT/CRIT
>>689921261
If your statement was true, then the crit chance would increase to 100% for the second hit, which contradicts the statement given in the question.
>>689921147
It actually doesn't work that way.
Your first hit is 50/50. If you crit then your next hit is 50/50. If you don't crit then your next hit is 100%. And there's also no possibility that a no crit scenario goes down.
Either the first hit is a crit, the second hit is a crit, or they're both crits. Each has the same probability of occurring. So it's 1/3.
That's obviously 1/3
>>689921667
It literally says it in the picture. One hit is always a crit. Irony much?
>>689921616
or for people who can't do math(s)
TT = 1/4
TH = 1/4
HT = 1/4
HH = 1/4
>>689921459
Why does the order even matter? One hit is critical who gives a fuck if it was the second or first hit? One is critical the other has. 50/50 chance of being critical too. 100%x50%=50%
>>689912733
So we're pitting statistics against probability here, that's why it's a famous thought experiment.
you can take the direct question and ignore the supporting information, .5*.5 = 25%
you can use the supporting information to remove an outcome, 25/75 = 33%
or since one of the outcomes is given you can calculate only one hit 1*.5 = 50%
>>689916188
in theory its possible but in reality its next to impossible due to the normal distribution curve
>>689921470
Thats one outcome that will happen 50% of the time. the other 2 will happen 25% of the time. YOU won't crit on the first shot 50% of the time. this isn't hard to understand
>>689921785
Just stop screaming and look it up instead.
>>689921535
Yes, but the first one doesn't have to be a crit. Which means that when you do the first hit and it's not a crit, you have already lost. No need to try the second hit. However, when your first hit does end up being a crit, you have a 50% of getting the second crit.
When you total all of the possibilities, only in one of those, you win. One out of three makes you win, which is ~33.33% winning chance.
>>689921763
What does that nigger know?
>>689921923
Why won't I crit on the first shot 50% of the time?
>>689921871
Thanks for the succinct appraisal of the situation.
>>689921833
Assuming your T means non-critical.
TT is not a possibility, since we know that "at least one of them is critical". True range is:
TH = 1/3
HT = 1/3
HH = 1/3
>>689921996
He knows that people like you should be liberal arts majors, for one.
>>689922052
You are. But if you don't crit on the first hit you HAVE to crit on the second one, that's the point.
>>689921551
you don't have 2 chance to roll crit, it's still 50%
>>689922277
Yes, but even if you hit that second crit, you lose. The point is getting both crits.
>>689921792
well the 33% contradicts the statement as well. Because you're calculations are skewed based on removing an outcome that isn't possible from the equation. Essentially what you guys are saying is
HH
TH
HT
You're saying you're going to crit your first hit 66% of the time.
>>689922277
crit on first, miss on second
crit on first, crit on second
miss on first, crit on second
three possibilities, one with 2 crits
1/3
whats the deal?
>>689913103
Done in 1.
One hit is already established.
So the odds of a second crit ride on 1 shot.
The odds of that shot being a crit are 50%
You have a penny on the table sitting heads side up. You are going to flip a penny. what are the odds that both show heads?
I though autists were supposed to be magic at math.
>>689922052
because your crit chance is 50%
>>689922432
No, the first one is 50% chance of getting it because the discarded option is still available for only one hit.
>>689922452
The first crit wasn't shot yet, you autist.
>>689912733
i'll just leave this here
https://en.wikipedia.org/wiki/Boy_or_Girl_paradox
>>689922432
and you're also going to crit you're second hit 66% of the time. Because 1 of the possibilities is having them both a crit.
>>689922439
I'm not the original fag you responded to, I just validated your implication that the chance of critting on the first hit is in fact 50% while giving additional information.